# Photon-likeness of hadron showers and impact of Lorentz boosting

### v1

@@ Abstract
We examine the probability of proton-induced air showers at $E>10$ EeV being misidentified as photon-induced due to neutral pions receiving a major part of the primary energy in the first interaction, thereby enhancing the electromagnetic shower component by their $\pi^{0}\to\gamma\gamma$ decay. Using CORSIKA simulations, we demonstrate the relevance of this effect at EeV energies. However, the probability for such photon-like events drops down strongly at the highest energies due to the increasing probability of Lorentz boosted $\pi^{0}$’s suffering hadronic interactions before decay. Different hadronic interaction models suggest that photon-like hadronic events may be observed at current UHECR observatories. A quantitative comparison of the observed number of background events found in recent photon searches published by the Pierre Auger Collaboration allows us conclude that the hypothesis of upwards fluctuations of $\pi^{0}$-production alone is insufficient to explain the data.
@@ 3 Photon-like event probability
The probability of finding large values of $k_{\Sigma\pi^{0}}$ depends primarily on the nature of the primary particle but is also found to depend on the hadronic interaction model used in the simulations. Both aspects are illustrated in Fig. 4. The probability for finding events with $k_{\Sigma\pi^{0}}\geq 0.7$ in simulations of proton primaries of $E_{0}>10$ EeV remains below 1‰, both for EPOS-LHC and QGSJetII-04. When comparing the two interaction models, we find the upwards fluctuations of $k_{\Sigma\pi^{0}}$ to be up to 10-20 times higher in EPOS-LHC as compared to QGSJetII-04. The impact of the primary particle type is even more significant: Only one event with $k_{\Sigma\pi^{0}}$ > 0.7 is found amongst $2\times 10^{5}$ simulated helium events while a few hundred events are observed for the same number of injected proton primaries, regardless of the hadronic interaction model.
@@ 4 Effect of the $\pi^{0}$ decay length at the highest energies
As discussed above, the observed probabilities of finding large values of $k_{\Sigma\pi^{0}}$ depends only very weakly on the primary energy. However, the impact that large values of $k_{\Sigma\pi^{0}}$ can have to the shower development is significant, once the energy of $\pi^{0}$’s exceed $\sim 1$ EeV. Above these energies, the $\pi^{0}$ decay length becomes sufficiently large to allow hadronic interactions taking place before the $\pi^{0}$ decays into two photons. For example, at $E_{\pi^{0}}=10$ EeV the decay length $\tau_{\rm decay}=\gamma\,c\,\tau_{0}$ (with $\gamma$ and $\tau_{0}$ being the Lorentz factor and lifetime of the $\pi^{0}$, respectively, and $c$ the speed of light) reaches about 2 km, which starts to exceed the pion-Air hadronic interaction length at atmospheric heights below about 6 km. Therefore, the enhancement of the electromagnetic component of hadron-induced air showers making them more photon-like due to upwards fluctuations of $k_{\Sigma\pi^{0}}$ will be counteracted at the highest energies by re-interactions of $\pi^{0}$ prior to their decay. The onset of $\pi^{0}$ re-interactions above cosmic ray energies of $\sim 1$ EeV has been noted also in d’Enterria et al. 2019.
@@ 5 Expectations for the irreducible background
It should be noted, on the other hand, that even modern air shower observatories, such as the Pierre Auger Observatory (Aab et al. 2015), will not achieve the ideal discrimination power between photons and hadrons shown in Fig. 8. Considering realistic air shower detection and reconstruction efficiencies, will inevitably increase the photon-like background, particularly when accounting for the much more abundant cosmic ray flux w.r.t. photons. It is beyond the scope of the present work to quantitatively analyze these effects. However, considering only the simulations, even hadron events with $k_{\text{boost}}$ as low as 0.6 would need to be classified incorrectly as photons, which seems improbable according to Fig. 8. Future studies may evaluate this discrimination capability for the background induced by protons.
@@ 6 Summary and conclusion
In this paper, the irreducible hadron background expected in searches for UHE photons at $E_{0}>10$ EeV has been studied and it was demonstrated that an upward fluctuation in the production of high energy $\pi^{0}$’s in the first cosmic ray interactions in the atmosphere can mimic the properties of photon induced air shower. This may either be due to a single high energy $\pi^{0}$ produced in the first interaction or due to several $\pi^{0}$’s carrying away a significant portion of the initial cosmic ray energy, both being quantified by the relative energy share $k_{\Sigma\pi^{0}}$. For a quantitative analysis, two hadronic interaction models were considered: EPOS-LHC and QGSJetII-04. It is found that the probability of finding events with large values of $k_{\Sigma\pi^{0}}$ is significantly higher for EPOS-LHC than for QGSJetII-04, so that EPOS is used for a conservative background estimate.
@@
By combining the calculated probabilities for finding different values of $k_{\text{boost}}$ with the actually measured number of cosmic ray events as a function of energy, we predict fewer photon-like events to be found in the data sample than are observed as background $N_{\text{bg}}$ in Abreu et al. 2023c for any reasonable selection of $k_{\text{boost}}$ ranges.
We note that the background estimates presented here are conservative in a number of respects so that theoretically expected hadron induced background is likely much lower, and could be increased only due to detector resolution and event reconstruction effects. However, based on our present understanding we conclude that the proton background caused by high energy $\pi^{0}$’s in the first interaction is unlikely to serve as an explanation for the identified photon candidates.

### v2

@@ Abstract
We examine the probability of proton-induced air showers at $E>10$ EeV being misidentified as photon-induced due to neutral pions receiving a major part of the primary energy in the first interaction, thereby enhancing the electromagnetic shower component by their $\pi^{0}\to\gamma\gamma$ decay. Using CORSIKA simulations, we demonstrate the relevance of this effect at EeV energies. However, the probability for such photon-like events drops down strongly at the highest energies due to the increasing probability of Lorentz boosted $\pi^{0}$’s suffering hadronic interactions before decay. Different hadronic interaction models suggest that photon-like hadronic events may be observed at current UHECR observatories. A quantitative comparison of the observed number of background events found in recent photon searches published by the Pierre Auger Collaboration suggests that the hypothesis of upward fluctuations in $\pi^{0}$-production alone is insufficient to explain the data. Detailed detector simulations will be needed to quantify the significance of these differences.
@@ 3 Photon-like event probability
The probability of finding large values of $k_{\Sigma\pi^{0}}$ depends primarily on the nature of the primary particle but is also found to depend on the hadronic interaction model used in the simulations. Both aspects are illustrated in Fig. 4. The probability for finding events with $k_{\Sigma\pi^{0}}\geq 0.7$ in simulations of proton primaries of $E_{0}>10$ EeV remains below 1‰, both for EPOS-LHC and QGSJetII-04. When comparing the two interaction models, we find the upward fluctuations of $k_{\Sigma\pi^{0}}$ to be up to 10-20 times higher in EPOS-LHC as compared to QGSJetII-04. The impact of the primary particle type is even more significant: Only one event with $k_{\Sigma\pi^{0}}$ > 0.7 is found amongst $2\times 10^{5}$ simulated helium events while a few hundred events are observed for the same number of injected proton primaries, regardless of the hadronic interaction model.
@@ 4 Effect of the $\pi^{0}$ decay length at the highest energies
As discussed above, the observed probabilities of finding large values of $k_{\Sigma\pi^{0}}$ depends only very weakly on the primary energy. However, the impact that large values of $k_{\Sigma\pi^{0}}$ can have to the shower development is significant, once the energy of $\pi^{0}$’s exceed $\sim 1$ EeV. Above these energies, the $\pi^{0}$ decay length becomes sufficiently large to allow hadronic interactions taking place before the $\pi^{0}$ decays into two photons. For example, at $E_{\pi^{0}}=10$ EeV the decay length $\tau_{\rm decay}=\gamma\,c\,\tau_{0}$ (with $\gamma$ and $\tau_{0}$ being the Lorentz factor and lifetime of the $\pi^{0}$, respectively, and $c$ the speed of light) reaches about 2 km, which starts to exceed the pion-Air hadronic interaction length at atmospheric heights below about 6 km. Therefore, the enhancement of the electromagnetic component of hadron-induced air showers making them more photon-like due to upward fluctuations of $k_{\Sigma\pi^{0}}$ will be counteracted at the highest energies by re-interactions of $\pi^{0}$ prior to their decay. The onset of $\pi^{0}$ re-interactions above cosmic ray energies of $\sim 1$ EeV has been noted also in d’Enterria et al. 2019.
@@ 5 Expectations for the irreducible background
It should be noted, on the other hand, that even modern air shower observatories, such as the Pierre Auger Observatory (Aab et al. (2015)), will not achieve the ideal discrimination power between photons and hadrons shown in Fig. 8. Considering realistic air shower detection and reconstruction efficiencies, will inevitably increase the photon-like background, particularly when accounting for the much more abundant cosmic ray flux w.r.t. photons. Moreover, the energy reconstruction ability of photon and photon-like air showers could play an additional role in estimating the background. It is beyond the scope of the present work to quantitatively analyze these effects. However, considering only the simulations, even hadron events with $k_{\text{boost}}$ as low as 0.6 would need to be classified incorrectly as photons, which seems improbable according to Fig. 8. Future studies may evaluate this discrimination capability for the background induced by protons.
@@ 6 Summary and conclusion
In this paper, the irreducible hadron background expected in searches for UHE photons at $E_{0}>10$ EeV has been studied and it was demonstrated that an upward fluctuation in the production of high energy $\pi^{0}$’s in the first cosmic ray interactions in the atmosphere can mimic the properties of photon induced air shower. This may either be due to a single high energy $\pi^{0}$ produced in the first interaction or due to several $\pi^{0}$’s carrying away a significant portion of the initial cosmic ray energy, both being quantified by the relative energy share $k_{\Sigma\pi^{0}}$. For a quantitative analysis, two hadronic interaction models were considered: EPOS-LHC and QGSJetII-04. It is found that the probability of finding events with large values of $k_{\Sigma\pi^{0}}$ is significantly higher for EPOS-LHC than for QGSJetII-04. In principle, more hadronic interaction models could be considered (e.g. Sibyll 2.3d (Riehn et al. (2020))). Newer models like the EPOS LHC-R (Pierog & Werner (2023)) could contribute to a more precise estimate of the irreducible proton background and might show more interesting features. However, for this study, EPOS-LHC can be used for a conservative background estimate.
@@
By combining the calculated probabilities for finding different values of $k_{\text{boost}}$ with the actually measured number of cosmic ray events as a function of energy, we find inconsistencies between the events found in the data sample and those observed as background $N_{\text{bg}}$ in Abreu et al. (2023c). The significance of this inconsistency depends on the benchmark $k_{\text{boost}}$, which needs to be determined by the discrimination power of the respective analysis.
We note that the background estimates presented here are conservative in a number of respects so that theoretically expected hadron induced background is likely much lower, and could be increased only due to detector resolution and event reconstruction effects. However, based on our present understanding we conclude that the proton background caused by high energy $\pi^{0}$’s in the first interaction is unlikely to serve as an explanation for the identified photon candidates. Nonetheless, integrating photon-like air shower simulations into photon searches (e.g. by the Pierre Auger Observatory) could yield valuable insights on how well the discrimination analysis performs. Vice versa, measured data could help to validate the findings of this study on the maximum irreducible proton background.